A 105 A current circulates around a 1.70-mm-diameter superconducting ring.

Part A

What is the ring's magnetic dipole moment?

Express your answer with the appropriate units.

2.38×10−4 A⋅m2

SubmitMy AnswersGive Up

Correct

Part B

What is the on-axis magnetic field strength 5.40 cmfrom the ring?

Express your answer with the appropriate units.

2.98•10−6T

SubmitMy AnswersGive Up

Incorrect; Try Again; 4 attempts remaining
Answer #1

given currrent is 105A

diammeter of the ring is 1.7mm

now we have to find the

What is the ring's magnetic dipole moment?

What is the on-axis magnetic field strength 5.40 cmfrom the ring?

1) The magnetic moment ? = I*A = 105A*?*r^2 =105*?*(0.85x10^-3m)^2 = 2.38x10^-4A-m^2


2) Bx = ?o*I*r^2/(2*(r^2 + z^2)^3/2 = 4?x10^-7*105*(0.85x10^-3)^2/(2*((0.85x10... + 0.054^2)^3/2 =

5.62x10^-6T